# WarmSwap: Sharing Dependencies for Accelerating Cold Starts in Serverless Functions

### v1

@@ 4.2 Comparing WarmSwap and Baseline
Figure 5a shows that WarmSwap outperforms the Baseline on cold-start latency for serverless functions that require large dependencies. WarmSwap is slightly slower than Baseline for the lightweight serverless functions (namely, helloworld, pyase, chameleon, and json_ dumps_load). However, WarmSwap significantly outperforms Baseline for serverless functions that need larger packages (i.e., “pandas”, “numpy”, “scikit-learn”, “torch”, and “keras”). WarmSwap is 1.2x, 1.8x, and 2.2x faster than Baseline for “lr _serving”, “cnn _serving”, and “rnn _serving”, respectively.
@@
Figure 5b shows the warm-start latency of WarmSwap and Baseline. For all six serverless functions, WarmSwap has similar warm-start latencies with Baseline.
@@
The cold start results show that WarmSwap has a performance advantage only on certain serverless functions. Specifically, WarmSwap is designed to solve the large dependency-loading problem in the serverless cold start. The lightweight functions only import standard and small Python packages. So directly booting those functions is faster than using WarmSwap. The lightweight functions incur additional overhead because WarmSwap needs to communicate with the dependency manager and migrate the dependency image into the containers, which increases some overhead.
@@
The warm start results show that WarmSwap’s performance is almost the same as that of the Baseline containers. During cold starts, the user code is already executed and necessary pages are already installed. WarmSwap does not add overhead to the warm start of the same function. Furthermore, there is no difference between the execution times for regular software dependencies and for migrated dependencies.
4.3 Components of Cold Start Execution Time
@@
To further analyze WarmSwap when executing a cold start, and answer RQ2, we decompose the cold-start time of WarmSwap and Baseline into three parts. Recall the architecture of WarmSwap from Figure 4a. Communication is the process by which the “migration client” communicates with the dependency manager and receives the process metadata. The process metadata is a special data structure associated with checkpoints that support CRIU’s lazy migration protocol (see Section 3.2). Migration is the process in which the dependency loader executes until the start of user-specific handler functions. The definitions of Execution, Boot, and Depedency Initialization are the same as those in Section 2.3.
@@
Figure 6: Cold-start latency breakdown for WarmSwap (H) and Baseline (B). The performance gain of WarmSwap arises because the migration is faster than the software dependency initialization of a Baseline container.
@@
Figure 6 shows the breakdown of cold-start latency for WarmSwap. The communication time is almost constant for all tested serverless functions, ranging from 25 ms to 94 ms according to the sizes of process metadata. Table 3 shows the sizes of process metadata for dependency images for each function. The size of process metadata is much smaller than the checkpoint image of the entire dependency. Since WarmSwap already migrates most of the memory by the page server, it does not need to transfer much additional information in the process metadata.
Compared with the dependency-loading time of Baseline, WarmSwap has a faster migration time in the case of functions that have larger dependencies. WarmSwap is 2.2x, 3.2x, and 2.5x faster than Baseline for booting dependencies for “lr _serving”, “cnn _serving” and “rnn _serving”, respectively. That explains how WarmSwap helps with the cold-start time. Since all memory pages are pre-initialized in the dependency image, user functions directly use them after migration. Within WarmSwap, the migration time dominates the cold start because migration is the key phase that transfers the dependency image that is needed by the user function.
4.4 Analysis of Benefits of WarmSwap Components
This section evaluates the contribution of the different optimization techniques used by WarmSwap in cold starts: live migration (RQ3) and different migration policies (RQ4). Since WarmSwap shows performance improvements only for functions with larger dependencies, we focus on the results for “lr_serving,” “cnn_serving,” and “rnn_serving” in this experiment.
Two migration policies of WarmSwap from Section 3.3 are evaluated:
WarmSwap Initiative Restore:
The page server copies all memory pages in parallel after the first “Page Fault.”
WarmSwap Lazy Restore:
The page server copies only the accessed pages that trigger a “Page Fault” at a time.
Two variants of WarmSwap are created as follows:
WarmSwap without Page Server:
WarmSwap without Lazy Migration:
Prototypes Cold Start (s) Warm Start (s)
WarmSwap Bulk Restore 2.67 2.06 0.89 1.80 0.65 0.004
WarmSwap Lazy Restore 2.93 3.29 0.76 2.39 2.71 0.004
w/o Page Server 3.00 2.14 0.91 1.82 0.68 0.004
w/o Lazy Migration 3.09 2.13 0.89 1.81 0.65 0.004
Table 2: Comparison of cold-start and warm-start results with: WarmSwap bulk restore, WarmSwap Lazy Restore; WarmSwap without Page Server; and WarmSwap without Lazy Migration.
Table 2 shows the cold-start and warm-start times for WarmSwap and its variants. WarmSwap with bulk restore has the best latency for “lr_serving” and “cnn_serving”, while WarmSwap with lazy restore outperforms the others for “rnn_serving”. The cold-start performance difference between bulk restore and lazy restore is related to the number of page faults a function triggers. The overhead of lazy restore increases significantly as the number of page faults rises. This is because whenever the function triggers a page fault, it needs to pause and wait for the page server to install the specific memory pages.
For WarmSwap without Page Server, we need to copy a relatively large image to the user container and restore all memory pages before execution. The overhead of copying the image is related to the checkpoint sizes, which are listed in Table 3.
For WarmSwap without Lazy Migration, the cold-start time is slower than WarmSwap because it needs to transfer all the memory pages through the network. Note that for “lr _serving”, WarmSwap without Lazy Migration is even slower than WarmSwap without page server. This is because “lr _serving” has a smaller checkpoint image size than the other functions.
All three prototypes (not counting Lazy Copy) have similar warm start times, which demonstrates that those three migration methods do not impact the time for warm starts. For Lazy Copy, the performance of warm starts decreases because “lr_serving” and “cnn_serving” trigger extra page faults during the first warm start. The performance will become normal for the following warm starts.
Based on these experiments, we believe bulk restore is more suitable in practice since the provider usually does not know how many page faults the target program will trigger. Moreover, it has no impact on warm-start performance.
Functions Metadata Size (MB) Image Size (MB)
@@ 4.4 Analysis of Benefits of WarmSwap Components
Table 3: The size of data that needs to be transferred from the worker node to the function container of WarmSwap and WarmSwap without Page Server.
@@
4.5 Case Study: Dependency Image Sharing Across Serverless Functions
In this section, we conduct simulation experiments to show the benefit of WarmSwap on sharing dependency images across different serverless functions to answer RQ5. We modify the “rnn_serving” functions by adding new code (using NumPy and torch) to the original user handler to create 10 different serverless functions. Assuming these 10 serverless functions are less frequently invoked, there is at most one instance for each function. However, the provider would still wish to gain the benefits of warm start even on these functions, if the cost is reasonable, in order to reduce the cold-start overhead.
Figure 7 compares WarmSwap and a checkpoint image-based method, Prebaking [23]. We implement Prebaking following the description in their paper, by checkpointing the entire previous warm serverless container with CRIU [7] to a function image and restoring it during the cold start. Note that, unlike WarmSwap and Baseline, the cost of container creation is larger in the case of Prebaking. Therefore, Figure 7 also includes the overhead of container creation for both methods.
Figure 7: Comparison of the cold-start latency and required cache/memory of WarmSwap and Prebaking Image. The experiment is based on 10 different serverless functions that require the same dependency image (Python+numpy+torch).
Figure 7 shows the accumulated cold-start latency (the sum of average cold-start latency for the 10 functions), and the required cache in memory needed by the optimization. WarmSwap significantly outperforms Prebaking in the cold-start performance (4.8 times faster), since restoring a large container with CRIU is much slower than migrating only the dependency.
@@ 4.5 Case Study: Dependency Image Sharing Across Serverless Functions
Furthermore, the provider only needs 260 MB memory to warm up the 10 different functions when using WarmSwap. In Prebaking, the provider needs to keep 10 checkpoint images (about 1.8 GB) in the cache for the 10 different functions. Hence, warming up the less frequently used functions with WarmSwap is cheaper than Prebaking.
@@
Although function-specific tuning methods, like vHive [25] and Catalyzer [8], achieve faster container restore, the provider must create and keep all function images in cache to optimize the less frequently used serverless functions. Thus, WarmSwap provides a practical method for providers to optimize cold starts for all functions at a reasonable cost.
5 Related Work
5.1 Function-Specific Tuning
One straightforward way to perform function-specific tuning is by using the Checkpoint/Restore techniques like CRIU [7]. One can create a checkpoint image of a specific warm serverless function and restore that image instead of initializing it from scratch during a cold start. For example, Prebaking [23] utilized transparent checkpoint technology (CRIU) to locally checkpoint the serverless function. This approach involves restarting the process during cold starts.
Catalyzer, as presented by Du et al. [8], confronts the intricate task of swiftly restoring serverless sandboxes, containers, virtual machines, and more. This innovative approach adeptly reinstates a virtualization-based function instance from a structured checkpoint image. Moreover, Catalyzer advances this restoration process by enabling on-demand recovery of both user-level memory states and system states. Notably, Catalyzer introduces the concepts of “sandbox fork” and “fork boot” to effectively tackle the challenge of warm starts, particularly when one or more hot instances already exist.
Building on checkpoint images, researchers have proposed more advanced optimizations. The authors of Catalyzer extended their contributions through MITOSIS [26]. This work takes the notion of sandbox forking further by leveraging Remote Direct Memory Access (RDMA) for remote forking. This innovation enables rapid warm starts for multiple sandboxes and supports stateful serverless computing. This extension enhances the versatility and applicability of the sandbox forking concept, facilitating efficient warm start strategies and accommodating stateful serverless scenarios.
Building upon the foundational work of Catalyzer, another study[25] introduces a noteworthy approach termed “hot page record-and-prefetch technology” aimed at expediting sandbox restoration and serverless function execution. The underlying concept revolves around recognizing that not all data within the C/R image will be utilized following restoration.
Additionally, the SEUSS framework [6] presents a unique C/R-like solution. In contrast to other approaches, SEUSS focuses on the restoration of function unikernel snapshots, sidestepping resource-intensive initialization steps. An intriguing aspect of SEUSS is its incorporation of page-level sharing across the entire software stack required for function execution. SEUSS encompasses not only unikernel binaries but also runtime snapshots and user code snapshots.
There are more ways to conduct function-specific tuning to speed up cold starts by analyzing user function code. Fasslight [17] optimizes code loading latency by loading only indispensable code. To implement this, Fasslight constructs the function-level call graph, separates other code, and loads only indispensable code to reduce cold start overhead.
Function-specific tuning usually achieves extremely high-speed cold starts, as reported by MITOSIS [26] and vHive [25]. However, as we analyzed in Section 2, these methods are not suitable for all serverless functions in production, especially those that are less frequently used. Additionally, some methods are difficult to implement on the cloud provider side because they require the provider to record [25] or analyze user code [17] or containers.
5.2 Infrastructure Optimization
For infrastructure optimization, the easiest way is to use a better hardware infrastructure. Further, there are plenty of system designs that help faster loading and initializing the functions by virtualization and caching systems.
Firecracker [1] by AWS Lambda is a Virtual Machine Monitor that utilizes the Linux Kernel’s KVM virtualization infrastructure to offer lightweight virtual machines. Firecracker significantly reduces the boot time of serverless instances and enhances switching performance.
AWS further reports its container optimization, which has implemented sharing across serverless functions in its approach to loading containers on-demand [5]. This is achieved through transforming serverless containers into data chunks, which are then deduplicated [18, 27] and shared across functions. These data chunks are efficiently stored within a three-tiered cache system. The primary focus of this work is on optimizing storage and facilitating the swift movement of larger serverless containers, all aimed at accelerating cold-start times.
The inherent nature of serverless functions has opened avenues for sharing not only within the same function but also across different functions. This is particularly valuable for frequently utilized functions that may accumulate millions of instances. Developers often opt for identical containers, runtimes, or even packages. In a recent report, AWS revealed that a significant 80% of newly uploaded serverless containers lack new data chunks [5].
Building on this insight, researchers extend infrastructure optimization through shareable resources. Pagurus [14], for instance, explores container-level sharing. Recognizing that certain user functions extensively share software packages, this approach allows new functions to borrow containers from existing ones, thereby reducing cold-start overhead. However, it’s crucial to acknowledge that sharing entire containers could introduce additional security challenges. HotC [24] focuses on reusing container runtimes to minimize container initialization overhead. HotC maintains a live container runtime pool, analyzes user input or configuration files, and provides available runtimes for immediate reuse.
In this work, we propose a feasible solution to extend infrastructure optimization for reducing dependency loading overhead.
5.3 Dependency Loading Overhead
Several previous works have recognized the importance of software dependency loading overhead. The SOCK framework, as explored in Oakes et al.’s work [21], analyzed the impact of loading Python packages on overhead. Their findings revealed that incorporating numerous popular libraries during the import process adds approximately 100 milliseconds to the cold-start time. Consequently, they devised an innovative strategy involving the creation of a lean container and a three-tiered Python library caching system. This approach effectively accelerates cold-start times.
Similarly, the FlashCube framework [16] addresses the delays caused by initializing function runtimes. To tackle this issue, FlashCube introduces a novel container framework. Instead of constructing a container from scratch, FlashCube assembles it by combining a selection of pre-existing general container components, including namespaces, cgroups, and runtimes. This approach maintains a template runtime process for each programming language (such as Python, Java, and Node.js). Photons [9] also recognize language runtime and library redundancy. They co-locate multiple instances of the same function within the same runtime, allowing concurrent invocations to share dependencies and application states.
Note that existing solutions attempt to pre-warm dependencies at the container level, requiring providers to pre-warm containers or container components for serverless functions. In contrast, WarmSwap is a lightweight solution that can be easily deployed across multiple nodes while effectively addressing the dependency loading problem.
6 Conclusion
In conclusion, this paper presents WarmSwap, a provider-side solution for accelerating dependency loading in serverless cold starts in production serverless platforms. WarmSwap is a complementary work that extends infrastructure optimization in serverless environments by representing software dependencies as middleware. It is designed to benefit the many less frequently used serverless functions, which is often not practical in a function-specific approach. WarmSwap is primarily beneficial when the dependencies of the serverless functions are large and when they incur significant latency during initialization. Containers with short initialization times already perform well during cold start, but recent years have seen greater use of complex software dependencies like AI/ML. It is in this regime that WarmSwap is expected to be used.
By using WarmSwap, a provider can implement a single pool of dependency images across its many computer hosts. This relieves the user of additional work. Since dependency images represent public and commonly used middleware, they can be shared among a whole category of serverless functions. Unlike previous research, the provider only needs to pre-warm and migrate the dependency process itself instead of entire containers. We implemented a prototype of WarmSwap based on offline AWS Lambda containers, and the evaluation shows that WarmSwap reduces the dependency loading overhead and warm-up cost across many serverless functions.

### v3

@@ IV-B Comparing HotSwap and Baseline
Figure 5a shows that HotSwap outperforms the Baseline on cold-start latency for serverless functions that require large dependencies. HotSwap is slightly slower than Baseline for the lightweight serverless functions (namely, helloworld, pyase, chameleon, and json_dumps_load). In production, HotSwap would be configured to avoid caching the trivial dependencies associated with these small serverless functions. Nevertheless, HotSwap significantly outperforms Baseline for serverless functions that need larger packages (i.e., “pandas”, “NumPy”, “scikit-learn”, “torch”, and “keras”). HotSwap is 1.2x, 1.8x, and 2.2x faster than Baseline for “lr _serving”, “cnn _serving”, and “rnn _serving”, respectively.
@@
Figure 5b shows the warm-start latency of HotSwap and Baseline. For all six serverless functions, HotSwap has similar warm-start latencies with Baseline.
@@
The cold start results show that HotSwap has a performance advantage only on certain serverless functions. Specifically, HotSwap is designed to solve the large dependency-loading problem in the serverless cold start. The lightweight functions only import standard and small Python packages. So directly booting those functions is faster than using HotSwap. The lightweight functions incur additional overhead because HotSwap needs to communicate with the dependency manager and migrate the dependency image into the containers, which increases some overhead.
@@
The warm start results show that HotSwap’s performance is almost the same as that of the Baseline containers. During warm starts, the user code is already executed, and necessary pages are already installed. HotSwap does not add overhead to the warm start of the same function. Furthermore, there is no difference between the execution times for regular software dependencies and for migrated dependencies.
IV-C Components of Cold Start Execution Time
@@
To analyze HotSwap’s cold-start performance and address RQ2, we decompose the cold-start time of HotSwap and Baseline into three parts. Recall the architecture of HotSwap from Figure 4a. Communication is the process by which the “migration client” communicates with the dependency manager and receives the process metadata. The process metadata is a special data structure associated with checkpoints that support CRIU’s lazy migration protocol (see Section III-C). Migration is the process in which the dependency loader executes until the start of the user-specific handler function. The definitions of Execution, Boot, and Dependency Initialization are the same as those in Section II-C.
@@
Figure 6: Cold-start latency breakdown for HotSwap (H) and Baseline (B). The performance gain of HotSwap arises because the migration is faster than the software dependency initialization from scratch.
@@
Figure 6 shows the breakdown of cold-start latency for HotSwap. The communication time is almost constant for all tested serverless functions, ranging from 25 ms to 94 ms according to the sizes of process metadata. Table II shows the sizes of process metadata for dependency images for each function. The size of process metadata is much smaller than the checkpoint image of the entire dependency.
TABLE II: The size of data that needs to be transferred from the worker node to the function container of HotSwap and HotSwap without Page Server.
Functions Metadata (MB) Image (MB)
@@
Compared with the dependency-loading time of Baseline, HotSwap has a faster migration time in the case of functions that have larger dependencies. HotSwap is 2.2x, 3.2x, and 2.5x faster than Baseline for booting dependencies for “lr _serving”, “cnn _serving” and “rnn _serving”, respectively. That explains how HotSwap helps with the cold-start time. Since all memory pages are pre-initialized in the dependency image, user functions directly use them after migration. Within HotSwap, the migration time dominates the cold start because migration is the key phase that transfers the dependency image that is needed by the user function.
IV-D Ablation Study
This section evaluates the contribution of the different optimization techniques used by HotSwap in cold starts: live migration (RQ3); and different migration policies (RQ4). Since HotSwap shows performance improvements only for functions with larger dependencies, we focus on the results for “lr_serving,” “cnn_serving,” and “rnn_serving”.
TABLE III: Comparison of cold-start and warm-start results with: HotSwap bulk restore, HotSwap Lazy Restore; HotSwap without Page Server; and HotSwap without Lazy Migration.
Prototypes Cold Start (s) Warm Start (s)
HotSwap Bulk Restore 2.67 2.06 0.89 1.80 0.65 0.004
HotSwap Lazy Restore 2.93 3.29 0.76 2.39 2.71 0.004
w/o Page Server 3.00 2.14 0.91 1.82 0.68 0.004
w/o Lazy Migration 3.09 2.13 0.89 1.81 0.65 0.004
Two migration policies of HotSwap from Section III-D are evaluated:
HotSwap Bulk Restore:
The page server copies all memory pages in parallel after the first “page fault”.
HotSwap Lazy Restore:
The page server copies only the accessed pages that trigger a “page fault”.
Two variants of HotSwap are created as follows:
HotSwap without Page Server:
HotSwap without Lazy Migration:
Table III shows the cold-start and warm-start latencies for HotSwap and its variants. HotSwap with bulk restore has the best latency for “lr_serving” and “cnn_serving”, while HotSwap with lazy restore outperforms the others for “rnn_serving”. The cold-start performance gap between bulk and lazy restore depends on page faults, as each fault in lazy restore pauses execution to wait for the page server.
For HotSwap without a Page Server, we copy a relatively large image to the user container and restore all memory pages before execution. The overhead of copying is related to the checkpoint sizes listed in Table II.
For HotSwap without Lazy Migration, the cold-start time is slower than HotSwap because it needs to transfer all the memory pages through the network before the start of function execution. Note that for “lr _serving”, HotSwap without Lazy Migration is even slower than HotSwap without page server. This is because “lr _serving” has a smaller checkpoint image size than the other functions.
All three prototypes (excluding Lazy Restore) have similar warm-start times, demonstrating that these migration methods have negligible impact on warm-start latency. For Lazy Restore, the performance of warm starts decreases because “lr_serving” and “cnn_serving” trigger extra page faults during the first warm start.
Based on these experiments, we believe bulk restore is more suitable in practice since the provider usually does not know how many page faults the target program will trigger.
@@ IV-D Ablation Study
IV-E Case Study: Dependency Image Sharing Across Serverless Functions
In this section, we conduct simulation experiments to demonstrate the benefits of HotSwap in sharing dependency images across different serverless functions, addressing RQ5. We use real-world serverless function traces from Azure [11], previously discussed in Section II-B of this paper, as the basis for our tests. We modified the “rnn_serving” function by adding new code (utilizing NumPy and Torch) to create 10 different serverless functions. For each function, we assigned random traces and measured latency over a two-week period. The keep-alive time was set to $15$ minutes. Since HotSwap is designed to optimize performance for less frequently used functions, we tested it under various invocation frequencies.
This experiment compares HotSwap with a checkpoint image-based method, Prebaking [13, 14]. Prebaking is chosen as a well-known system that employs CRIU-based checkpoint/restore to accelerate serverless cold starts. We implemented Prebaking as described in their paper, by checkpointing the entire previously warm serverless container using CRIU [27] to create a function image, which is then restored during a cold start. To enhance fairness, we store all prebaking images in memory. Note that the cold start latency measured in this experiment includes the container creation overhead for both methods.
Figure 7: Comparison of average latency and required cache/memory between HotSwap and Prebaking[13, 14] under Azure Function traces [11]. The experiment involves 10 different serverless functions, all using the same dependency image ( NumPy + Torch). Each function is assigned a real-world invocation trace from Azure. The left figure shows latency across different invocation rate groups, ranging from the lowest to the highest quartile.
Figure 7 shows the average latency and required cache for HotSwap and Prebaking. The average latency is calculated from all invocations of the ten different functions. To evaluate different invocation rate scenarios, we categorized the Azure traces into quartiles based on invocation rate. The “Lowest Quartile” represents experiments on function invocation traces with rates below the 25th percentile, while the other three groups correspond to “25-50%”, “50-75%”, and above the 75th percentile, respectively.
The latency results indicate that HotSwap outperforms Prebaking by 1.4x across all trace groups. This improvement is due to HotSwap’s superior cold-start performance. The average latency decreases as the invocation rate increases, resulting in more warm starts. Notably, the 75th percentile of invocation rates from Azure traces [11] is 0.04 calls per minute. This demonstrates that the majority of serverless functions are, in fact, used infrequently.
@@ IV-E Case Study: Dependency Image Sharing Across Serverless Functions
Furthermore, the provider requires only 260 MB of memory to warm up the 10 different functions when using HotSwap. In contrast, Prebaking requires the provider to maintain 10 checkpoint images (approximately 2.4 GB) in the cache for the same functions. Therefore, warming up the less frequently used functions with HotSwap is more cost-effective than Prebaking. This benefit can increase further if there are more functions that share the same dependency image.
@@
Although function-specific tuning methods, like vHive [17] and Catalyzer [15], achieve faster container restore, the provider must create and keep all function images in cache to optimize the less frequently used serverless functions. Thus, HotSwap provides a practical method for providers to optimize cold starts for skewed workloads at a reasonable cost.
V Related Work
Pre-Warm. One straightforward way to optimize cold starts is to prevent them from happening. Prior research has explored various approaches such as adjusting keep-alive policies, leveraging application-level composition [28], mining invocation patterns [9], and applying machine learning models [29, 30, 31], in order to predict function usage and proactively pre-warm containers [10, 8, 11]. However, these strategies often lead to resource inefficiency due to inaccurate predictions, especially in dynamic production environments. HotSwap takes a different approach by directly reducing cold-start overhead instead of relying on speculative pre-warming.
Function-Specific Optimization. A common method for reducing cold-start latency is leveraging Checkpoint/Restore (C/R) techniques such as CRIU [27]. Prebaking [13, 14] applies CRIU to checkpoint entire serverless functions, significantly speeding up start-up across runtimes. Catalyzer [15] improves sandbox restoration by efficiently resuming serverless functions across sandboxes. MITOSIS [18] extends this approach by introducing remote sandbox forking using RDMA.
Further extending this idea, Ustiugov et al. [17] propose a “hot page record-and-prefetch” technique to accelerate sandbox restoration and function execution. SEUSS [16] takes a unikernel-based approach, restoring function snapshots while avoiding resource-heavy initialization. Fasslight [32] optimizes code loading latency by loading only essential code. Although effective for high-frequency serverless functions, these methods may not suit skewed workloads in production, where checkpointing and analysis overhead can outweigh the benefits. Furthermore, a recent user-perspective study [33] suggests that Node.js performs better than Java for infrequently invoked functions. However, providers have no control over which language users choose to implement their functions.
Cross-Function Optimization and Dependency Loading. Cross-function optimizations improve performance by sharing resources across functions. Pagurus [34] reduces cold-start overhead via container reuse, while HotC [35] maintains a pool of preloaded runtimes for reuse based on function inputs.
Software dependency loading is a key factor in cold-start latency. SOCK [19] uses pre-initialized Zygote containers and multi-tier caching to speed up Python library loading. (In fact, SOCK uses a generalization of the original Android-based Zygotes.) FlashCube [21] builds containers from prebuilt components to reduce initialization time. Photons [36] co-locates function instances in a shared runtime to enable dependency reuse. TrEnv [20] leverages remote memory pools via CXL/RDMA to share execution environments. SLIMSTART [7] removes infrequently used libraries through runtime profiling and automated transformation. HotSwap aims to introduce a more lightweight and easily implementable solution to reduce dependency-related overhead.
Infrastructure and Storage Systems. Infrastructure optimization can further reduce cold-start latency by improving storage and caching mechanisms. Firecracker [23], a micro-VM developed by AWS Lambda, leverages the Linux Kernel’s KVM virtualization to provide lightweight virtual machines, significantly reducing boot times and improving context-switching performance. AWS also optimizes serverless container sharing by deduplicating data chunks across functions, revealing that 80% of uploaded containers contain redundant data [22, 37, 38]. These infrastructure optimizations are complementary to HotSwap, as they focus on different aspects of serverless performance improvements.
VI Limitations
HotSwap is specifically designed for to be active during the dependency loading phase. HotSwap targets the overhead associated with dependency loading and is most effective when this overhead is substantial. For functions with only trivial dependencies, using HotSwap to accelerate dependency loading may not be necessary. From the provider’s perspective, this also means that it is not required to cache all dependencies, only those that are popular and large.
Evaluation on Other Runtimes and Languages. In this paper, we evaluate HotSwap only on the Python runtime. We chose Python because it is one of the most popular languages for serverless functions, particularly in data processing, automation, and machine learning tasks. Moreover, many widely used Python libraries are large and contribute significantly to cold start overhead. However, since HotSwap relies on process-level migration, it could, in theory, be applied to other runtimes and languages as well—especially in scenarios where dependency loading overhead is a major concern.
Centralized Dependency Management. The Dependency Manager introduced in Section III-B is centralized in our prototype design and may become a bottleneck when handling numerous migration requests. In practice, cloud providers can implement HotSwap as a distributed caching system with multiple Dependency Managers operating within a cluster.
VII Conclusion
In conclusion, this paper presents HotSwap, a cross-function optimization designed to accelerate dependency loading in serverless cold starts. HotSwap provides a strong benefit in cases of skewed workloads, with many infrequently used serverless functions (the problem of a “long tail” in the distribution). It primarily optimizes the dependency loading phase, a major bottleneck in current production clouds. By enabling a shared pool of dependency images across multiple compute hosts, HotSwap allows providers to efficiently manage commonly used middleware. We implemented a HotSwap prototype based on offline AWS Lambda containers, and our evaluation demonstrates that it significantly reduces dependency loading overhead and warm-up costs across various serverless functions.
Acknowledgment. We thank our shepherd, Yehia Elkhatib, and the anonymous reviewers for their constructive feedback. This work is supported in part by NSF Awards 1910601 and 2124897.